OpenStudy (anonymous):

if sina+sinb=-21/65,cosa+cosb=-27/65 then cos(a-b/2)=?

OpenStudy (rajee_sam):

Sin a + Sin b = 2 Sin (a + b)/2 . Cos (a-b)/2 Cos a + Cos b = 2 Cos (a + b)/2. Cos ( a-b)/2 square them and add them 4 sin² ( a + b)/2 . Cos² (a-b) /2 + 4 Cos² (a + b)/2 . Cos² (a- b)/2 4 Cos² (a-b)/2 [ Sin² (a + b)/2 + Cos² (a + b)/2] = 4 Cos² (a-b)/2 . 1 since we squared and added. We have to square the values and add them too. (-21/65)² + (-27/65)² = 4 Cos² (a-b)/2 (1170/4225) = 4 Cos ² (a-b)/2 now you can find the value by doing the last few steps right??

OpenStudy (rajee_sam):

@vanimanasa1998

OpenStudy (anonymous):

thank u...